# Training PINNs with Hard Constraints and Adaptive Weights: An Ablation Study

### v2

@@ 2.2 Mini-batching
We have eliminated the IC/BC losses, but the transformed PDE loss might create additional difficulties in training due to the introduction of IC/BC functions into the PDE. We adopt a machine learning technique, called mini-batching, commonly used to reduce training complexity by evaluating the loss on randomly selected subsets of collocation data. In PINNs, this reduces memory costs, especially when handling large numbers of collocation points. Sampling mini-batches from collocation points also helps stabilize and accelerate optimization. So we consider mini-batching in the training for decreasing the PDE residual loss. As pointed out in wight2020solving , the mini-batching technique is similar to time-marching resampling. But with mini-batching, one does not employ an explicit time grid.
2.3 Self-Adaptive Physics-Informed Neural Networks
@@
The proper choice of penalty term $\lambda$ can be updated from derived formula as mentioned in wang2020and . However, data-driven discovery of $\lambda$ is possible. Outline in mcclenny2020self , the loss is further modified as
$\text{Data Loss}(\lambda_{IC,i}\}_{i=1}^{N_{IC}},\{\lambda_{BC,i}\}_{i=1}^{N_{BC}})=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}m(\lambda_{IC,i})|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
$\quad+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}m(\lambda_{BC,i})|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2}$
@@
Physical $\text{Physical Loss}(\{\lambda_{PDE,i}\}_{i=1}^{N_{PDE}})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}m(\lambda_{PDE,i})|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}$
@@
The set of loss weights, $\{\lambda_{IC,i}\}_{i=1}^{N_{IC}}\cup\{\lambda_{BC,i}\}_{i=1}^{N_{BC}}\cup\{\lambda_{PDE,i}\}_{i=1}^{N_{PDE}}$, are designed in a way to increase when needed during any training epoch in order to combat the spectral bias. Adaptive methods are essential to ensure that the neural network accurately addresses the challenging spots in the solutions of “stiff” PDEs. In our experiments, we employ self-adaptive PINNs as proposed by mcclenny2020self to train PINNs adaptively. This approach involves fully trainable adaptation weights applied individually to each training point, allowing the neural network to independently identify and focus on the challenging regions of the solution.
2.4 Performance Measures
@@
$\mathcal{E}_{1}=\frac{\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|}{\sum_{i=1}^{N}|u(x_{i},t_{i})|},\quad\quad\quad\mathcal{E}_{2}=\frac{\Big{(}\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|^{2}\Big{)}^{1/2}}{\sqrt{\sum_{i=1}^{N}|u(x_{i},t_{i})|^{2}}}.$
These metrics represent the normalized differences between the model predictions and the ground truth values, providing a quantitative evaluation of model performance.
3 Numerical Experiments
In this section, we examine the proposed PINN scheme on a diverse set of prototypical stiff time-dependent PDEs with periodic boundary conditions, aiming to demonstrate its robustness and flexibility in handling a broad class of challenging equations. These benchmarks exhibit a wide range of physical and mathematical characteristics: from stiff dynamics (e.g., Kuramoto–Sivashinsky), to multi-species reaction–diffusion systems (e.g., Gray–Scott and Belousov–Zhabotinsky), and nonlinear dispersive waves (e.g., nonlinear Schrödinger equation). Such equations often involve strong nonlinearities, stiffness, and parameter sensitivities, posing significant challenges for data-driven methods such as PINNs. Nonetheless, our results illustrate the capacity of the proposed framework to overcome these difficulties and accurately approximate the solutions across these varied regimes. Detailed experimental results are presented in Section 3.1.
To begin, we employ Latin Hypercube Sampling to generate collocation points across the full spatio-temporal domain, including its boundary. To facilitate training and better incorporate the initial condition into the solution structure, we adopt the following transformation:
@@ 3 Numerical Experiments
$\tilde{u}(t,\mathbf{x})=\exp(-Ct)*u_{0}(\mathbf{x})+t*u_{nn}(t,\mathbf{x}),\quad(t,\mathbf{x})\in[0,T]\times\bar{\Omega};\quad\psi=\exp(-Ct)*u_{0}(\mathbf{x}),\,\phi=t.$
where $u_{0}(\mathbf{x})$ denotes the initial condition, and we test two values of $C\in\{0.1,1\}$, to assess how the decay rate affects learned dynamics. The transformed PINN $u_{nn}$ is trained to minimize a loss function that accounts for the PDE residuals, which now includes IC/BC. To rigorously assess the accuracy of our results, we use Chebfun, a spectral-style computing system in MATLAB that enables highly accurate function approximations. This allows us to make precise comparisons between the ground-truth solutions and the PINN-generated solutions.
For each benchmark, we not only present the main numerical results but also carry out detailed ablation studies to evaluate the contribution of each proposed technique. In each ablation experiment, we deactivate one component of the methodology at a time, keeping the remaining elements active under the same hyper-parameter configurations, to measure the relative $L^{2}$ error and relative $L^{1}$ error. The neural network architecture used throughout consists of 7 hidden layers with 32 neurons per layer and tanh activation functions. The initial weights are fixed across all ablation runs for consistency. Model training proceeds in two stages: first using the Adam optimizer kingma2014adam with a learning rate of $10^{-3}$ for $50,000$ steps, followed by fine-tuning via the L-BFGS-B optimizer with a learning rate of $1$ for $5,000$ steps. Each experiment uses $16,384$ interior collocation points, along with $128$ initial and $128$ boundary data points sampled via LHS. For PDE systems (e.g., Gray–Scott), we further apply a mini-batching strategy with $5$ batches and a batch size of $3,280$ to enhance training stability.
Overall, these extensive experiments highlight the effectiveness and adaptability of our PINN scheme in solving a wide variety of stiff time-dependent PDEs. The results provide not only strong empirical validation but also a reliable and reproducible benchmark for future developments in this area. For full experimental details and visualizations, see Section 3.1.
3.1 Allen-Cahn Equation $I$

### v3

@@ A.1 Mini-batching
We have eliminated the IC/BC losses, but the transformed PDE loss might create additional difficulties in training due to the introduction of IC/BC functions into the PDE. We adopt a machine learning technique, called mini-batching, commonly used to reduce training complexity by evaluating the loss on randomly selected subsets of collocation data. In PINNs, this reduces memory costs, especially when handling large numbers of collocation points. Sampling mini-batches from collocation points also helps stabilize and accelerate optimization. So we consider mini-batching in the training for the PDE residual loss. As pointed out in (Wight & Zhao, 2020), the mini-batching technique is similar to time-marching resampling but without an explicit time grid.
A.2 Self-Adaptive PINNs
@@
The proper choice of penalty term $\lambda$ can be updated from derived formula as mentioned in (Wang et al., 2020b). However, data-driven discovery of $\lambda$ is possible. Outline in (McClenny & Braga-Neto, 2020), the loss is further modified as
$\text{Data Loss}(\{\lambda^{IC}_{i}\}_{i=1}^{N_{IC}},\{\lambda^{BC}_{i}\}_{i=1}^{N_{BC}})$
$=\frac{1}{N_{IC}}\sum_{i=1}^{N_{IC}}F_{mask}(\lambda^{IC}_{i})|u_{nn}(0,\mathbf{x}_{i}^{IC})-u_{0}(\mathbf{x}_{i}^{IC})|^{2}$
$+\sum_{i=1}^{d}\frac{1}{N_{BC}}\sum_{j=1}^{N_{BC}}F_{mask}(\lambda^{BC}_{j})|u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC})-u_{nn}(t_{j}^{BC},\mathbf{x}_{j}^{BC}+P\mathbf{e}_{i})|^{2},$
and
@@
$\text{Physical Loss}(\{\lambda^{CL}_{i}\}_{i=1}^{N_{CL}})=\frac{1}{N_{CL}}\sum_{i=1}^{N_{CL}}F_{mask}(\lambda^{CL}_{i})|(\frac{\partial u}{\partial t}-\mathcal{P}(u_{nn})-f)(t_{i}^{CL},\mathbf{x}_{i}^{CL})|^{2}.$
@@
The set of loss weights, $\{\lambda^{IC}_{i}\}_{i=1}^{N_{IC}}\cup\{\lambda^{BC}_{i}\}_{i=1}^{N_{BC}}\cup\{\lambda^{CL}_{i}\}_{i=1}^{N_{CL}}$, are designed in a way to increase when needed during any training epoch in order to combat the spectral bias. Adaptive methods are essential to ensure that the neural network accurately addresses the challenging spots in the solutions of “stiff” PDEs. In our experiments, we employ self-adaptive PINNs as proposed by (McClenny & Braga-Neto, 2020) to train PINNs adaptively. This approach involves fully trainable adaptation weights applied individually to each training point, allowing the neural network to independently identify and focus on the challenging regions of the solution.
A.3 Performance Measures
@@
$\small\mathcal{E}_{1}=\frac{\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|}{\sum_{i=1}^{N}|u(x_{i},t_{i})|},\quad\mathcal{E}_{2}=\frac{\sqrt{\sum_{i=1}^{N}|\mathcal{U}(x_{i},t_{i})-u(x_{i},t_{i})|^{2}}}{\sqrt{\sum_{i=1}^{N}|u(x_{i},t_{i})|^{2}}}.$
These metrics measure the discrepancy between the predicted and reference solutions in average (energy-based) and absolute (mass-based) senses, respectively. The relative form normalizes the error with respect to the reference solution, enabling fair comparisons across different scales and variables. To rigorously assess the accuracy of our results, we implemented a semi-implicit spectral method in Matlab. This allows us to make precise comparisons between the ground-truth solutions $u(x_{i},t_{i})$ and the PINN-generated solutions $\mathcal{U}(x_{i},t_{i})$.
A.4 Proofs of the Theorems
We present the proofs for the two aforementioned theorem.
Theorem (Well Conditioning).
The transformation is well defined when $\phi$ and $\psi$ is $C^{1}$ in time and have $K^{th}$ order partial derivatives w.r.t $\mathbf{x}$. Moreover, $\phi_{t}(0,\mathbf{x})\neq 0$. Then $u_{nn}$ will satisfy a new PDE
@@ A.4 Proofs of the Theorems
$\begin{cases}&\phi\partial_{t}u_{nn}+\mathcal{P}[\psi+\phi u_{nn}]+\phi_{t}u_{nn}=f-\psi_{t},\quad(t,\mathbf{x})\in(0,T]\times\Omega,\\ &u_{nn}(0,\mathbf{x})=\frac{u_{t}(0,\mathbf{x})-\psi_{t}(0,\mathbf{x})}{\phi_{t}(0,\mathbf{x})},\quad\mathbf{x}\in\bar{\Omega}.\end{cases}$
Proof.
Since $\mathcal{L}_{s}[u]=\partial_{t}u+\mathcal{P}[u]$ is applied to the transformation $u=\psi+\phi u_{nn}$ and $u_{nn}$ is constructed with hyper tangent as the activation function, we only need $\psi$ and $\phi$ to satisfy the required regularity from $\mathcal{L}_{s}$ so that the operator can still be applied, hence $\psi$ and $\phi$ have to have first-order time derivative and $K^{th}$-order partial derivatives w.r.t to $\mathbf{x}$, we further require them to be continuous in order to allow classical solutions to exist for $\mathcal{L}_{s}$. Meanwhile from the transformation we have
$u_{nn}(t,\mathbf{x})=\frac{u(t,\mathbf{x})-\psi(t,\mathbf{x})}{\phi(t,\mathbf{x})},\quad\text{for $(t,\mathbf{x})\in(0,T]\times\Omega$.}$
As $t\rightarrow 0$, we have $\psi(0,\mathbf{x})=u_{0}(\mathbf{x})$ and $\phi(0,\mathbf{x})=0$, hence by L’Hôpital’s rule
$\lim_{t\rightarrow 0}u_{nn}(t,\mathbf{x})=\lim_{t\rightarrow 0}\frac{u(t,\mathbf{x})-\psi(t,\mathbf{x})}{\phi(t,\mathbf{x})}=\lim_{t\rightarrow 0}\frac{u_{t}(t,\mathbf{x})-\psi_{t}(t,\mathbf{x})}{\phi_{t}(t,\mathbf{x})},$
when $\phi_{t}(0,\mathbf{x})\neq 0$, then limit is defined and we can have
$u_{nn}(0,\mathbf{x})=\lim_{t\rightarrow 0}u_{nn}(t,\mathbf{x})=\frac{u_{t}(0,\mathbf{x})-\psi_{t}(0,\mathbf{x})}{\phi_{t}(0,\mathbf{x})}.$
Moreover
$\mathcal{L}_{s}(\psi+\phi u_{nn})=\partial_{t}(\psi+\phi u_{nn})+\mathcal{P}[\psi+\phi u_{nn}]=\phi\partial_{t}u_{nn}+\mathcal{P}[\psi+\phi u_{nn}]+\phi_{t}u_{nn}+\psi_{t},$
hence the new PDE for $u_{nn}$. ∎
Next, the theorem on Hessian bounds.
Theorem.
Assume that $\mathcal{P}$ is linear and $\phi(t,\mathbf{x})=\phi(t)\neq 0$ for $t>0$, then for a one-hidden layer network, we have for a network $u(\mathbf{y};\theta)$ with $\bm{\theta}\in\mathbb{R}^{D}$ being the network parameters, the following relationship is satisfied
$u(\mathbf{y};\bm{\theta})=\mathbf{h}^{\top}(\mathbf{y})\bm{\theta},\quad\mathbf{h}=\begin{bmatrix}h_{1}(\mathbf{y})&\cdots&\mathbf{h}_{D}(\mathbf{y})\end{bmatrix}^{\top}$,
where $\mathbf{y}=\mathbf{z}=(t,\mathbf{x})\in(0,T]\times\Omega$ or $\mathbf{y}=\bm{\zeta}=(0,\mathbf{x}^{IC})\in\{0\}\times\bar{\Omega}$. Therefore the two losses become
$\text{Loss}_{s}(u(\mathbf{y};\bm{\theta}))$ $=\frac{1}{2N}|\mathbf{A}\bm{\theta}-\mathbf{f}|_{\ell^{2}(\mathbb{R}^{N})}^{2}+\frac{1}{2M}|\mathbf{C}\bm{\theta}-\mathbf{u}_{0}|_{\ell^{2}(\mathbb{R}^{M})}^{2},\quad\text{(soft constrained)}$
$\text{Loss}_{h}(u(\mathbf{z};\bm{\theta}))$ $=\frac{1}{2N}|\bm{\Lambda}_{\phi}\mathbf{A}\bm{\theta}+\bm{\Lambda}_{\phi_{t}}\mathbf{B}\bm{\theta}-\tilde{\mathbf{f}}|_{\ell^{2}(\mathbb{R}^{N})}^{2},\quad\text{(hard constrained)}$
where
$\mathbf{A}=\begin{bmatrix}\mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{1})]\\ \vdots\\ \mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{N})]\end{bmatrix},\quad\mathbf{B}=\begin{bmatrix}\mathbf{h}^{\top}(\mathbf{z}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\mathbf{z}_{N})\end{bmatrix},\quad\text{and}\quad\mathbf{C}=\begin{bmatrix}\mathbf{h}^{\top}(\bm{\zeta}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\bm{\zeta}_{M})\end{bmatrix}$,
where $\mathbf{x}^{IC}_{i}\in\bar{\Omega}$; and for the diagonal matrices $\bm{\Lambda}_{\phi}$ and $\bm{\Lambda}_{\phi_{t}}$
$\bm{\Lambda}_{\phi}=\text{diag}(\phi(\mathbf{z}_{1}),\cdots,\phi(\mathbf{z}_{N}))\quad\text{and}\quad\bm{\Lambda}_{\phi_{t}}=\text{diag}(\phi_{t}(\mathbf{z}_{1}),\cdots,\phi_{t}(\mathbf{z}_{N})),$
and the vectors
$\mathbf{f}=\begin{bmatrix}f(\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})\end{bmatrix},\quad\tilde{\mathbf{f}}=\begin{bmatrix}f(\mathbf{z}_{1})-\mathcal{L}_{s}[\psi](\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})-\mathcal{L}_{s}[\psi](\mathbf{z}_{N})\end{bmatrix},\quad\text{and}\quad\mathbf{u}_{0}=\begin{bmatrix}u_{0}(\mathbf{x}_{1}^{IC})\\ \vdots\\ u_{0}(\mathbf{x}_{M}^{IC})\end{bmatrix}$.
Therefore the Hessians of the losses are
$H_{s}(\bm{\theta})$ $=\frac{1}{N}\mathbf{A}^{\top}\mathbf{A}+\frac{1}{M}\mathbf{C}^{\top}\mathbf{C},$
$H_{h}(\bm{\theta})$ $=\frac{1}{N}(\mathbf{A}^{\top}\bm{\Lambda}_{\phi}^{2}\mathbf{A}+2\mathbf{A}^{\top}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\phi_{t}}\mathbf{B}+\mathbf{B}^{\top}\bm{\Lambda}_{\phi_{t}}^{2}\mathbf{B}),$
Then the Hessians are bounded as
$|H_{s}|_{2}$ $\leq\frac{|\mathcal{L}_{s}|_{2}^{2}}{N}|\mathbf{B}|_{2}^{2}+\frac{1}{M}|\mathbf{C}|_{2}^{2},$
$|H_{h}|_{2}$ $\leq\frac{1}{N}(C_{1}|\mathcal{L}_{s}|_{2}^{2}+2C_{2}|\mathcal{L}_{s}|_{2}+C_{3})|\mathbf{B}|_{2}^{2}.$
where $|\mathcal{L}_{s}|_{2}$ is the $L^{2}$-operator norm of $\mathcal{L}_{s}$ and
$C_{1}=\max_{\mathbf{z}}\phi^{2}(\mathbf{z}),\quad C_{2}=\max_{\mathbf{z}}|\phi(\mathbf{z})\phi_{t}(\mathbf{z})|,\quad\text{and}\quad C_{3}=\max_{\mathbf{z}}\phi_{t}^{2}(\mathbf{z}).$
Proof.
Once we have $u(\mathbf{y};\bm{\theta})=\mathbf{h}^{\top}(\mathbf{y})\bm{\theta}$, then the loss from the soft-constrained PINN can be re-written as
$\text{Loss}_{s}(u(\bm{\theta}))=\frac{1}{2N}\sum_{i=1}^{N}|(\mathcal{L}_{s}[u(\cdot;\bm{\theta})]-f)(\mathbf{z}_{i})|^{2}+\frac{1}{2M}\sum_{j=1}^{M}|u(\bm{\zeta}_{i};\bm{\theta})-u_{0}(\mathbf{x}_{i}^{IC})|^{2},$
If we define the matrices
$\mathbf{A}=\begin{bmatrix}\mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{1})]\\ \vdots\\ \mathcal{L}_{s}[\mathbf{h}^{\top}(\mathbf{z}_{N})]\end{bmatrix}\quad\text{and}\quad\mathbf{C}=\begin{bmatrix}\mathbf{h}^{\top}(\bm{\zeta}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\bm{\zeta}_{M})\end{bmatrix}$,
and the vectors
$\mathbf{f}=\begin{bmatrix}f(\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})\end{bmatrix}\quad\text{and}\quad\mathbf{u}_{0}=\begin{bmatrix}u_{0}(\mathbf{x}_{1}^{IC})\\ \vdots\\ u_{0}(\mathbf{x}_{M}^{IC})\end{bmatrix}$.
then $\text{Loss}_{s}(u(\bm{\theta}))=\frac{1}{2N}|\mathbf{A}\bm{\theta}-\mathbf{f}|_{\ell^{2}(\mathbb{R}^{N})}^{2}+\frac{1}{2M}|\mathbf{C}\bm{\theta}-\mathbf{u}_{0}|_{\ell^{2}(\mathbb{R}^{M})}^{2}$. Similarly, for the loss
$\text{Loss}_{h}(u(\bm{\theta}))=\frac{1}{2N}\sum_{i=1}^{N}|(\mathcal{L}_{s}[\psi(\cdot)+\phi(\cdot)u(\cdot;\bm{\theta})]-f)(\mathbf{z}_{i})|^{2},$
Since
$\mathcal{L}_{s}[\psi+\phi u]=\phi(u_{t}+\mathcal{P}[u])+\phi_{t}u+\psi_{t},\quad\text{$\mathcal{P}$ is linear and $\phi(t,\mathbf{x})=\phi(t)$}.$
Hence with
$\mathbf{B}=\begin{bmatrix}\mathbf{h}^{\top}(\mathbf{z}_{1})\\ \vdots\\ \mathbf{h}^{\top}(\mathbf{z}_{N})\end{bmatrix}\quad\text{and}\quad\tilde{\mathbf{f}}=\begin{bmatrix}f(\mathbf{z}_{1})-\mathcal{L}_{s}[\psi](\mathbf{z}_{1})\\ \vdots\\ f(\mathbf{z}_{N})-\mathcal{L}_{s}[\psi](\mathbf{z}_{N})\end{bmatrix}$,
and the two diagonal matrices
$\bm{\Lambda}_{\phi}=\text{diag}(\phi(\mathbf{z}_{1}),\cdots,\phi(\mathbf{z}_{N}))\quad\text{and}\quad\bm{\Lambda}_{\phi_{t}}=\text{diag}(\phi_{t}(\mathbf{z}_{1}),\cdots,\phi_{t}(\mathbf{z}_{N})),$
we have $\text{Loss}_{h}(u(\bm{\theta}))=\frac{1}{2N}|\bm{\Lambda}_{\phi}\mathbf{A}\bm{\theta}+\bm{\Lambda}_{\phi_{t}}\mathbf{B}\bm{\theta}-\tilde{\mathbf{f}}|_{\ell^{2}(\mathbb{R}^{N})}^{2}$. Then the two gradients are
$\nabla_{\bm{\theta}}\text{Loss}_{s}(u(\bm{\theta}))$ $=\frac{1}{N}(\mathbf{A}^{\top}\mathbf{A}\bm{\theta}-\mathbf{A}^{\top}\mathbf{f})+\frac{1}{M}(\mathbf{C}^{\top}\mathbf{C}\bm{\theta}-\mathbf{C}^{\top}\mathbf{u}_{0}),$
[…]
Differentiating w.r.t to $\bm{\theta}$ again, we obtain the Hessians as
$H_{s}(\bm{\theta})$ $=\frac{1}{N}\mathbf{A}^{\top}\mathbf{A}+\frac{1}{M}\mathbf{C}^{\top}\mathbf{C},$
$H_{h}(\bm{\theta})$ $=\frac{1}{N}(\mathbf{A}^{\top}\bm{\Lambda}_{\phi}^{2}\mathbf{A}+2\mathbf{A}^{\top}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\phi_{t}}\mathbf{B}+\mathbf{B}^{\top}\bm{\Lambda}_{\phi_{t}}^{2}\mathbf{B}),$
Recall that $\mathcal{L}_{s}[u]=u_{t}+\mathcal{P}[u]$, hence by the property of an $L^{2}$ function/operator norm on $\mathcal{L}_{s}$, we have
$|\mathcal{L}_{s}[u]|_{2}\leq|\mathcal{L}_{s}|_{2}|u|_{2},\quad\text{This $|\cdot|_{2}$ is a norm on functions.}$
Then we have $|\mathbf{A}^{\top}\mathbf{A}|_{2}\leq|\mathcal{L}_{s}|_{2}^{2}|\mathbf{B}|_{2}^{2}$ (Here $|\mathbf{A}|_{2}$ is a matrix norm). And if we define
$C_{1}=\max_{\mathbf{z}}\phi^{2}(\mathbf{z}),\quad C_{2}=\max_{\mathbf{z}}|\phi(\mathbf{z})\phi_{t}(\mathbf{z})|,\quad\text{and}\quad C_{3}=\max_{\mathbf{z}}\phi_{t}^{2}(\mathbf{z}).$
We can have
$|\mathbf{A}^{\top}\bm{\Lambda}_{\phi}^{2}\mathbf{A}|_{2}$ $\leq C_{1}|\mathbf{A}^{\top}\mathbf{A}|_{2}\leq|\mathcal{L}_{s}|_{2}^{2}|\mathbf{B}|_{2}^{2},$
$|\mathbf{A}^{\top}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\phi_{t}}\mathbf{A}|_{2}$ $\leq C_{2}|\mathbf{A}^{\top}\mathbf{B}|\leq C_{2}|\mathcal{L}_{s}|_{2}|\mathbf{B}|_{2}^{2},$
$|\mathbf{B}^{\top}\bm{\Lambda}_{\phi_{t}}\mathbf{B}|_{2}$ $\leq C_{3}|\mathbf{B}^{\top}\mathbf{B}|_{2}\leq C_{3}|\mathbf{B}|_{2}^{2}$
Putting them back in, we obtain the final bounds
$|H_{s}|_{2}$ $\leq\frac{|\mathcal{L}_{s}|_{2}^{2}}{N}|\mathbf{B}|_{2}^{2}+\frac{1}{M}|\mathbf{C}|_{2}^{2},$
$|H_{h}|_{2}$ $\leq\frac{1}{N}(C_{1}|\mathcal{L}_{s}|_{2}^{2}+2C_{2}|\mathcal{L}_{s}|_{2}+C_{3})|\mathbf{B}|_{2}^{2}.$
Appendix B Examples: Additional Details
We provide additional details for the ablation study, comparison study, $2D$ application and sensitivity study of $m$ here. We also explain why each PDE or PDE system is chosen due to their particular difficulties for solvers in each sub-section.
B.1 Implementation Details
All experiments were implemented in PyTorch ($v2.2$) and executed on both CUDA-enabled NVIDIA GPUs and Apple Silicon devices via the MPS backend. Our code runs efficiently across Linux (Ubuntu $22.04$), macOS (Sonoma), and Windows $11$ platforms. Primary training was conducted on an NVIDIA $A100$ GPU ($40$ GB memory) and a MacBook Pro with Apple M1 Pro chip ($16$-core GPU, $32$ GB unified memory).
For each benchmark, we used a fully connected neural network with $7$ hidden layers, each containing $32$ neurons and tanh activation functions, to represent the latent solution functions. The same initial weights were used across all ablation experiments for consistency.
Model training was carried out using a two-stage strategy: first, we trained using the Adam optimizer with an initial learning rate of $10^{-3}$ for $50,000$ steps. This was followed by fine-tuning with the L-BFGS-B optimizer using a learning rate of $1$ for $5,000$ steps. We did not perform hyperparameter sweeps, and the total number of training epochs was fixed across all experiments.
This paper does not rely on any datasets. We used Latin Hypercube Sampling (LHS) to generate a fixed set of $16,384$ interior collocation points, along with $128$ initial data points and $128$ boundary data points. This same sampling configuration was used consistently across all experiments to ensure comparability. For PDE systems, we employed a mini-batching strategy with $5$ batches of size $3,280$ to improve memory efficiency during training.
Ablation studies were conducted by disabling one component of our method at a time, while keeping all other settings fixed, to evaluate its contribution using relative $L^{2}$ and relative $L^{1}$ error metrics, which measure the discrepancy between the predicted and reference solutions in average (energy-based) and absolute (mass-based) senses, respectively. The relative form normalizes the error with respect to the reference solution, enabling fair comparisons across different scales and variables.
We also compare HC-PINN with causal training (MLP) and enhanced RBA using the same data and a total of 50k iterations, instead of the 300k iterations used in the original works. For the sensitivity study, experiments with different $m$ values are conducted on the same points, using the same number of iterations and the same architecture. For 2D training, we use 4,0964{,}096 initial collocation points, 10,00010{,}000 residual collocation points per segment, and $800$ boundary points per segment. The architecture is the same as in the 1D training.
B.2 Allen-Cahn PDE